# PHENIX Collaboration

### v1

@@ Abstract
The PHENIX experiment measured the centrality dependence of two-pion Bose-Einstein correlation functions in $\sqrt{s_{{}_{NN}}}=200$ GeV Au$+$Au collisions at the Relativistic Heavy Ion Collider at Brookhaven National Laboratory. The data are well represented by Lévy-stable source distributions. The extracted source parameters are the correlation-strength parameter $\lambda$, the Lévy index of stability $\alpha$, and the Lévy-scale parameter $R$ as a function of transverse mass $m_{T}$ and centrality. The $\lambda(m_{T})$ parameter is constant at larger values of $m_{T}$, but decreases as $m_{T}$ decreases. The Lévy scale parameter $R(m_{T})$ decreases with $m_{T}$ and exhibits proportionality to the length scale of the nuclear overlap region. The Lévy exponent $\alpha(m_{T})$ is independent of $m_{T}$ within uncertainties in each investigated centrality bin, but shows a clear centrality dependence. At all centralities, the Lévy exponent $\alpha$ is significantly different from that of Gaussian ($\alpha=2$) or Cauchy ($\alpha=1$) source distributions. Comparisons to the predictions of Monte-Carlo simulations of resonance-decay chains show that in all but the most peripheral centrality class (50%–60%), the obtained results are inconsistent with the measurements, unless a significant reduction of the in-medium mass of the $\eta^{\prime}$ meson is included. In each centrality class, the best value of the in-medium $\eta^{\prime}$ mass is compared to the mass of the $\eta$ meson, as well as to several theoretical predictions that consider restoration of $U_{A}(1)$ symmetry in hot hadronic matter.
@@ I Introduction
In a previous paper [1] on Bose-Einstein Correlations (BECs)—also known as Hanbury Brown-Twiss (HBT) Correlations—the PHENIX Collaboration found that for 0%–30% centrality Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$=200 GeV, the two-particle BECs are well-described by a Lévy-stable source distribution. The traditional description using a Gaussian source distribution is inadequate. A strong preference for the Lévy description had also been seen in e++e-collisions at the Large Electron-Positron Collider [2] and in $p$$+$$p$, $p$$+$Pb , and Pb$+$Pb collisions at the Large Hadron Collider [3, 4, 5, 6] and in Be+Be collisions at the Super Proton Synchrotron [7].
@@
Presented here is a precise measurement of the centrality and transverse-mass dependence of the two-pion BEC function in Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$= 200 GeV by the PHENIX experiment at the Relativistic Heavy Ion Collider (RHIC). This data sample, recorded in 2010, allows a fine transverse-mass binning and inference of the shape of the correlation function more precisely than was possible with earlier data sets. The source parameters of the Lévy distribution ($\lambda$, $R$, $\alpha$)  [^1] Note that followed are the conventions introduced in Ref. [8, 9, 10], which is based on a book by P. J. Nolan [11] on univariate Lévy-stable source distributions, including also multivariate, but symmetric Lévy-stable distributions. are measured, as was done for the 0%–30% centrality class in Ref. [1]. For the first time, these results are presented as a function of centrality for six subsequent centrality classes in the range 0%–60%. The centrality and the transverse-mass dependence of the Lévy-fit parameters are characterized with simple, theoretically and empirically motivated fit functions. The centrality dependence of the parameters of these functions is investigated in detail.
@@
The structure of this paper is as follows: Sections II and III present the PHENIX experimental setup and the selection of the data sample, respectively. Section IV explores the procedure of measurement and fitting of the two-pion correlation function. Section V discusses the systematic uncertainties. Section VI presents the extracted Lévy parameters of the source as a function of the transverse mass of the pair for the six centrality intervals. Finally, Sections VII and VIII discuss the simulations and some of the possible physics interpretations of these results, which have a strong exclusion power due to their high precision. The discussions and Appendix employ Monte-Carlo simulations to interpret the PHENIX data.
In particular, PHENIX data on the transverse-mass and centrality dependence of the Lévy intercept parameter $\lambda(m_{T})$ are compared to centrality dependent Monte-Carlo simulations of resonance decay chains. In all but the most peripheral centrality class (50%–60%), the Monte-Carlo simulations are found to be inconsistent with the measurements unless a significant reduction of the in-medium mass of the $\eta^{\prime}$ meson is included. In each centrality class, the best value of the in-medium $\eta^{\prime}$ mass is determined from $\chi^{2}$ and confidence level (CL or p-value) maps, based on a comparison of PHENIX data and Monte-Carlo simulations. The resulting values of in-medium modified $\eta^{\prime}$ masses are compared to the mass of the $\eta$ meson as well as to several theoretical predictions that consider restoration of $U_{A}(1)$ symmetry in hot hadronic matter as discussed in Sections VII and VIII and then further detailed in the Appendix. Throughout this paper, units are used such that $\hbar=c=1$.
@@ III Data sample
The present analysis shares almost all details with the previous analysis of Ref. [1]. The similarities and differences between the current and previous analyses are detailed below. In the present analysis, the Lévy fit parameters are determined in 23 bins of transverse mass from 0.248 GeV to 0.876 GeV and in six, 10% wide centrality bins in the range of 0%–60%. Well measured tracks are selected using the same single-track cuts as in Ref. [1]. The event selection criteria (except the centrality selection) and the particle identification techniques are also the same as in Ref. [1]. The single-track cuts and their variations are considered as sources of systematic uncertainties, as described in Section V.
@@
These pair cuts were carefully investigated in the previous Lévy analysis [1]. However, due to the different centrality selection, the pair cut settings here are slightly modified. The pair cuts are defined in the $\Delta\varphi$$\Delta z$ plane as:
$\Delta\varphi>\Delta\varphi_{0}\left(1-\frac{\Delta z}{\Delta z_{0}}\right)\hskip 5.69046pt{\rm and}\hskip 5.69046pt\Delta\varphi>\Delta\varphi_{1}\hskip 5.69046pt{(\rm DCandPbSc)},$ (1)
$\Delta\varphi>\Delta\varphi_{0}\left(1-\frac{\Delta z}{\Delta z_{0}}\right)\hskip 5.69046pt{(\rm TOFeast)},$ (2)
$\Delta\varphi>\Delta\varphi_{0}\hskip 5.69046pt{\rm and}\hskip 5.69046pt\Delta z>\Delta z_{0}\hskip 5.69046pt{(\rm TOFwest)}.$ (3)
@@
Table 1: The values of the coordinates for the pair cuts and the alternative values used to determine systematic uncertainties.
@@
Default cut settings .12 8. .017 .12 12 .075 14.0 .12 16 .015
Loose drift chamber cut .11 7. .016 .12 12 .075 14.0 .12 16 .015
Strict drift chamber cut .13 9. .018 .12 12 .075 14.0 .12 16 .015
Loose ID detector cuts .12 8. .017 .11 11 .070 13.0 .11 15 .013
Strict ID detector cuts .12 8. .017 .13 13 .080 15.0 .13 17 .017
@@ IV.1 Measuring the correlation function
In principle, a detailed shape analysis of the two-particle correlation functions could require three-dimensional measurements, but the lack of statistical precision could make such measurements impractical, because the goal here is to obtain precise results in several transverse mass and centrality bins. Thus, the correlation functions use a single variable $Q$. PHENIX preliminary results on a multivariate Lévy analysis are also available [13], but go beyond the scope of the present paper.
@@
which depends on the relative momentum only through the one-dimensional variable $Q$, the magnitude of the relative momentum of the pair in the LCMS. It was shown in Ref. [17] that the same quadratic sum appears in three-dimensional symmetric Lévy-stable distributions. So our choice for the relative momentum variable $Q$ actually includes, as a special case, the Lévy analysis of three-dimensional Bose-Einstein correlation functions under the condition that the HBT radii in all three spatial dimensions are equal within experimental uncertainties. An approximate equality of the transverse-mass and centrality dependent BEC radii was found in several experiments, e.g., S$+$Pb collisions at the CERN Super-Proton-Synchrotron energies by the NA44 Collaboration [18], and in $\sqrt{s_{NN}}=200$ GeV Au$+$Au collisions at $\sqrt{s_{NN}}=200$ GeV by the PHENIX and STAR Collaborations in the Gaussian approximation in Refs. [19, 20]. PHENIX preliminary multivariate Lévy fits in $\sqrt{s_{NN}}=200$ GeV Au$+$Au collisions are also used in Ref. [13]. Especially note that the PHENIX preliminary three-dimensional Lévy-analysis for Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$= 200 GeV finds the three BEC radii in the LCMS are approximately equal (except perhaps at small transverse mass).
@@
Our choice of $Q$ can be defined using the measured momenta of identified pions, $p_{i}^{\mu}=(E_{i},p_{i,x},p_{i,y},p_{i,z})$. For $i=1,2$,
@@
where the z axis coincides with the beam axis. The correlation function is measured as:
@@
In choosing the background distribution $B(Q)$, each member of the pairs are selected from different events. Hence, $B(Q)$ contains only trivial kinematic correlations between independent particles that are distributed with the same single-particle spectra and in the same kinematic range. The background distribution is affected by the centrality selection cuts and detector acceptance effects, rapidity, and transverse-mass cuts. In $A(Q)$ the pairs are correlated not only due to Bose-Einstein correlations, but also by other initial- or final-state interactions. Such effects from interactions include branching processes of jets, hadronization effects, resonance decays, energy and momentum conservation laws, possible other kinematic effects, such as elastic scattering, Coulomb, and strong final-state interactions.
@@
In general, $A(Q|MC)$ and $B(Q|MC)$ should be generated from Monte-Carlo simulations that describe these single-particle data and all other nonBEC correlation effects. These Monte-Carlo distributions are then used to correct $B(Q)$ for its lack of nonBEC correlations, as was done, for example in e++e- collisions by the L3 collaboration [2]. However, in high-energy heavy-ion physics there is a good reason why this complicated and Monte-Carlo-dependent procedure is not necessary. Namely, the presence of high-multiplicity events, with $\langle n\rangle\gg 1$, where $\langle n\rangle$ is the mean charged-particle multiplicity at midrapidity. The usual kinematic correlations, which are due to resonance decays or conservation laws are proportional to the mean multiplicity, $\langle n\rangle$, while Bose-Einstein correlations grow with the mean number of pairs $\langle n(n-1)\rangle$. Thus in high-multiplicity heavy ion collisions, Bose-Einstein correlations (together with Coulomb and strong final-state interactions that also grow proportionally with the number of pairs) outnumber all the other correlations. Hence, the MC approach can be safely abandoned. In the expression $A(Q)/B(Q)$ the Bose-Einstein, Coulomb, and strong final-state correlations dominate, and the other correlations are suppressed as $\langle n\rangle/\langle n(n-1)\rangle\propto 1/\langle n\rangle\ll 1$.
@@
In BEC measurements using charged particles, Coulomb repulsion modifies the correlation function at low $Q$ values creating the “Coulomb hole” [14]. To account for the final-state Coulomb interaction, the Coulomb wave function is integrated over the source of pions. Strong final-state interactions could also be taken into account using phase-shifts that modify the Coulomb wave function. However, the strong final-state interaction of the pion pairs is small compared to the experimental precision [14, 22]. All the Lévy fits here (without corrections for the strong final-state interactions) are of good quality with CL $\gg 0.1$%. The strong final-state interactions are found not to affect the quality of Lévy fits, nor to change the parameters significantly [14, 22]. Therefore, such corrections are not considered in this analysis. In contrast, the effect of Coulomb interactions is clearly visible as a Coulomb hole at small values of $Q$. The fitting function is appropriately modified, as detailed Section IV.2.2.
@@ IV.2.1 The Lévy shape
For the shape of the correlation function, it it not possible to $a$ $priori$ assume or know, what is or what should be the shape-model of the Bose-Einstein correlation functions. Measurements of this quantity test the hypothesis of being consistent with a Lévy shape. Using the plane-wave approximation and assuming a spherically symmetric, Lévy-type source, the two-particle BEC function has the simple form
@@
where $\lambda$ is the strength of the correlation, $R$ is the scale parameter in physical units, and $\alpha$ is the Lévy index of stability [17]. This hypothesis is tested not only by successful fits to the PHENIX data with CL $\gg 0.1$%, which are stable and robust with systematic variations of particle identification, acceptance and pair cuts, as detailed in Section V. The validity of the Lévy shape is also checked by employing a Lévy expansion technique [23, 9, 10]. This method utilizes a complete set of polynomials that are orthonormal with respect to a Lévy weight; so this method is able to characterize and model independently any deviation from a Lévy shape. To first order of the expansion, no significant deviation from the Lévy shape is found in any of the centrality and transverse-mass ranges investigated here.
The relevance of stable distributions to the analysis of (Coulomb-corrected) Bose-Einstein correlation was studied in Ref. [17], following the general mathematical ideas summarized in Ref. [11]. Univariate stable distributions are usually characterized by the Fourier transform of their density distributions that are called the characteristic functions. Following the convention of the previous analysis [1] and the theoretical paper where the idea first appeared [17], the $S(\alpha,\beta,\gamma,\delta;1)$ notation is used. The parameter $\alpha$ is the Lévy index of stability (or characteristic exponent) that is limited to the domain $0<\alpha\leq 2$. The asymmetry parameter $\beta$ is limited to the domain $-1\leq\beta\leq 1$. The scale parameter $\gamma$ is nonnegative and the location parameter $\delta$ can be any real number with $-\infty<\delta<\infty$. The book by P. J. Nolan [11] details exhaustively the $\alpha=1$ special case and the ubiquitous nature of Lévy-stable source distributions. It provides illustrations of the stable densities in the $S(\alpha,\beta,\gamma,\delta;1)$ parameterization.
@@
The correlation function is thus based on the assumption of the simplest case of univariate and symmetric ($\beta=0$) stable distributions. The scale parameter $\gamma$ is replaced, by introducing the physical parameter $R$ [17]. In the standardized notation of Nolan [11], this corresponds to the $(\alpha,\beta{=}0,\gamma{=}R/2^{\frac{1}{\alpha}},\delta;1)$ convention. The Fourier-transformed source-density distribution has a simple form, $f(q)=\exp(iq\delta-\frac{1}{2}|qR|^{\alpha})$ and its modulus square leads to the simple form of the BEC function of Eq. (8). Note that the correlation function does not depend on $\delta$.

### v2

@@ Abstract
The PHENIX experiment measured the centrality dependence of two-pion Bose-Einstein correlation functions in $\sqrt{s_{{}_{NN}}}=200$ GeV Au$+$Au collisions at the Relativistic Heavy Ion Collider at Brookhaven National Laboratory. The data are well represented by Lévy-stable source distributions. The extracted source parameters are the correlation-strength parameter $\lambda$, the Lévy index of stability $\alpha$, and the Lévy-scale parameter $R$ as a function of transverse mass $m_{T}$ and centrality. The $\lambda(m_{T})$ parameter is constant at larger values of $m_{T}$, but decreases as $m_{T}$ decreases. The Lévy scale parameter $R(m_{T})$ decreases with $m_{T}$ and exhibits proportionality to the length scale of the nuclear overlap region. The Lévy exponent $\alpha(m_{T})$ is independent of $m_{T}$ within uncertainties in each investigated centrality bin, but shows a clear centrality dependence. At all centralities, the Lévy exponent $\alpha$ is significantly different from that of Gaussian ($\alpha=2$) or Cauchy ($\alpha=1$) source distributions. Comparisons to the predictions of Monte-Carlo simulations of resonance-decay chains show that in all but the most-peripheral centrality class (50%–60%), the obtained results are inconsistent with the measurements, unless a significant reduction of the in-medium mass of the $\eta^{\prime}$ meson is included. In each centrality class, the best value of the in-medium $\eta^{\prime}$ mass is compared to the mass of the $\eta$ meson, as well as to several theoretical predictions that consider restoration of $U_{A}(1)$ symmetry in hot hadronic matter.
@@ I Introduction
In a previous paper [1] on Bose-Einstein Correlations (BECs)—also known as Hanbury Brown and Twiss (HBT) Correlations—the PHENIX Collaboration found that for 0%–30% centrality Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$=200 GeV, the two-particle BECs are well-described by a Lévy-stable source distribution. However, the traditional description of the same dataset, using a Gaussian source distribution was found to be inadequate [1]. A strong preference for the Lévy description had also been seen in e++e-collisions at the Large Electron-Positron Collider [2] and in $p$$+$$p$, $p$$+$Pb , and Pb$+$Pb collisions at the Large Hadron Collider [3, 4, 5, 6], in Be+Be and Ar+Sc collisions at the Super Proton Synchrotron [7, 8] and in Au$+$Au collisions at the Relativistic Heavy Ion Collider (RHIC) [9].
@@
Presented here is a precise measurement of the centrality and transverse-mass dependence of the two-pion BEC function in Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$= 200 GeV by the PHENIX experiment at RHIC. This data sample, recorded in 2010, allows a fine transverse-mass binning and inference of the shape of the correlation function more precisely than was possible with earlier data sets. For the first time, these results are presented as a function of centrality for six centrality classes in the range 0%–60%. As was done for the 0%–30% centrality class in Ref. [1], the source parameters of the Lévy distribution ($\lambda$, $R$, $\alpha$) are measured. [^1] Note that followed are the conventions introduced in Refs. [10] and  [11, 12, 13], which is based on a book by P. J. Nolan [14] on univariate Lévy-stable source distributions, including also multivariate, but symmetric Lévy-stable distributions.. The centrality and the transverse-mass dependence of the Lévy-fit parameters are characterized with simple, theoretically and empirically motivated fit functions. The centrality dependence of the parameters of these functions is investigated in detail.
@@
The structure of this paper is as follows: Sections II and III present the PHENIX experimental setup and the selection of the data sample, respectively. Section IV explores the procedure of measurement and fitting of the two-pion correlation function. Section V discusses the systematic uncertainties. Section VI presents the extracted Lévy parameters of the source as a function of centrality. Section VII discusses Monte Carlo simulations and some of the possible physics interpretations of these results, which have a strong exclusion power due to their high precision. Section VIII summarizes and concludes. Finally an Appendix details our Monte-Carlo simulations to interpret the PHENIX data.
[…]
@@ III Data sample
The present analysis shares almost all details with the previous analysis of Ref. [1], including using the word “cuts” to refer to selection criteria. The similarities and differences between the current and previous analyses are detailed below. In the present analysis, the Lévy fit parameters are determined in 23 bins of transverse mass from 0.248 GeV to 0.876 GeV and in six, 10% wide centrality bins in the range of 0%–60%. Well-measured tracks are selected using the same single-track cuts as in Ref. [1]. The event-selection criteria (except the centrality selection) and the particle-identification techniques are also the same as in Ref. [1]. The single-track cuts and their variations are considered as sources of systematic uncertainties, as described in Section V.
@@
These pair cuts were carefully investigated in the previous Lévy analysis [1]. However, due to the different centrality selections, the pair-cut settings here are slightly modified. The pair cuts are defined in the $\Delta\varphi$$\Delta z$ plane as:
[…]
@@
Table 1: The values of the coordinates for the pair-selection (cuts) criteria and the alternative values used to determine systematic uncertainties.
@@
Default cut settings 0.12 8. 0.017 0.12 12 0.075 14.0 0.12 16 0.015
Loose drift chamber cut 0.11 7. 0.016 0.12 12 0.075 14.0 0.12 16 0.015
Strict drift chamber cut 0.13 9. 0.018 0.12 12 0.075 14.0 0.12 16 0.015
Loose ID detector cuts 0.12 8. 0.017 0.11 11 0.070 13.0 0.11 15 0.013
Strict ID detector cuts 0.12 8. 0.017 0.13 13 0.080 15.0 0.13 17 0.017
@@ IV.1 Measuring the correlation function
In principle, a detailed shape analysis of the two-particle correlation functions could require three-dimensional measurements, but the lack of statistical precision could make such measurements impractical. The goal here is to obtain precise results in several transverse-mass and centrality bins. Thus, the correlation functions use a single variable $Q$. PHENIX preliminary results on a multivariate Lévy analysis are also available [16], but go beyond the scope of the present paper.
@@
which depends on the relative momentum only through the one-dimensional variable $Q$, the magnitude of the relative momentum of the pair in the LCMS. It was shown in Ref. [10] that the same quadratic sum appears in three-dimensional symmetric Lévy-stable distributions. So our choice for the relative momentum variable $Q$ actually includes, as a special case, the Lévy analysis of three-dimensional Bose-Einstein correlation functions under the condition that the HBT radii in all three spatial dimensions are equal within experimental uncertainties. An approximate equality of the transverse-mass and centrality-dependent BEC radii was found in several experiments, e.g., S$+$Pb collisions at the CERN Super-Proton-Synchrotron energies by the NA44 Collaboration [20], and in $\sqrt{s_{NN}}=200$ GeV Au$+$Au collisions at $\sqrt{s_{NN}}=200$ GeV by the PHENIX and STAR Collaborations in the Gaussian approximation in Refs. [21, 22]. PHENIX preliminary multivariate Lévy fits in Au$+$Au collisions at $\sqrt{s_{NN}}=200$ GeV are also used in Ref. [16]. Especially note that the PHENIX preliminary three-dimensional Lévy-analysis for Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$=200 GeV finds that the three BEC radii in the LCMS are approximately equal (except perhaps at small transverse mass).
@@
Our choice of $Q$ can be written using the measured momenta of identified pions, $p_{i}^{\mu}=(E_{i},p_{i,x},p_{i,y},p_{i,z})$. For $i=1,2$:
@@
where the z axis coincides with the beam axis.
The correlation function is measured as:
@@
In choosing the background distribution $B(Q)$, each member of the pairs are selected from different events. Hence, $B(Q)$ contains only trivial kinematic correlations between independent particles that are distributed with the same single-particle spectra and in the same kinematic range. The background distribution is affected by the centrality-selection cuts and detector-acceptance effects, rapidity, and transverse-mass cuts. In $A(Q)$ the pairs are correlated not only due to Bose-Einstein correlations, but also by other initial- or final-state interactions. Such effects from interactions include branching processes of jets, hadronization effects, resonance decays, energy, and momentum conservation laws, possible other kinematic effects, such as elastic scattering, Coulomb, and strong final-state interactions.
@@
In general, $A(Q|MC)$ and $B(Q|MC)$ should be generated from Monte-Carlo simulations that describe these single-particle data and all other nonBEC correlation effects. These Monte-Carlo distributions are then used to correct $B(Q)$ for its lack of nonBEC correlations, as was done, for example in e++e- collisions by the L3 collaboration [2]. However, in high-energy heavy-ion physics there is a good reason why this complicated and Monte-Carlo-dependent procedure is not necessary. Namely, the presence of high-multiplicity events, with $\langle n\rangle\gg 1$, where $\langle n\rangle$ is the mean charged-particle multiplicity at midrapidity. The usual kinematic correlations, which are due to resonance decays or conservation laws, are proportional to the mean multiplicity, $\langle n\rangle$, while Bose-Einstein correlations grow with the mean number of pairs $\langle n(n-1)\rangle$. Thus in high-multiplicity heavy ion collisions, Bose-Einstein correlations (together with Coulomb and strong final-state interactions that also grow proportionally with the number of pairs) outnumber all the other correlations. Hence, the MC approach can be safely abandoned. In the expression $A(Q)/B(Q)$ the Bose-Einstein, Coulomb, and strong final-state correlations dominate, and the other correlations are suppressed as $\langle n\rangle/\langle n(n-1)\rangle\propto 1/\langle n\rangle\ll 1$.
@@
In BEC measurements using charged particles, Coulomb repulsion modifies the correlation function at low $Q$ values creating the “Coulomb hole” [17]. To account for the final-state Coulomb interaction, the Coulomb wave function is integrated over the source of pions. Strong final-state interactions could also be taken into account using phase shifts that modify the Coulomb wave function. However, the strong final-state interaction of the pion pairs is small compared to the experimental precision [17, 24]. All the Lévy fits here (without corrections for the strong final-state interactions) are of good quality with CL $\gg 0.1$%. The strong final-state interactions are found not to affect the quality of Lévy fits, nor to change the parameters significantly [17, 24]. Therefore, such corrections are not considered in this analysis. In contrast, the effect of Coulomb interactions is clearly visible as a Coulomb hole at small values of $Q$. The fitting function is appropriately modified, as detailed Section IV.2.2.
@@ IV.2.1 The Lévy shape
For the shape of the correlation function, it is not possible to $a$ $priori$ assume or know, what is or what should be the shape-model of the Bose-Einstein correlation functions. Measurements of this quantity test the hypothesis of being consistent with a Lévy shape. Using the plane-wave approximation and assuming a spherically symmetric, Lévy-type source, the two-particle BEC function has the simple form
@@
where $\lambda$ is the strength of the correlation, $R$ is the scale parameter in physical units, and $\alpha$ is the Lévy index of stability [10]. This hypothesis is tested not only by successful fits to the PHENIX data with CL $\gg 0.1$%, which are stable and robust with systematic variations of particle identification, acceptance, and pair cuts, as detailed in Section V. The validity of the Lévy shape is also checked by employing a Lévy expansion technique [25, 12, 13]. This method utilizes a complete set of polynomials that are orthonormal with respect to a Lévy weight function [12]; so this method is able to characterize and model independently any deviation from a Lévy-stable source shape. To first order of the expansion, no significant deviation from the Lévy shape is found in any of the centrality and transverse-mass ranges investigated here.
The relevance of stable distributions to the analysis of (Coulomb-corrected) Bose-Einstein correlations was studied in Ref. [10], following the general mathematical ideas summarized in Ref. [14]. Univariate stable distributions are usually characterized by the Fourier transform of their density distributions that are called the characteristic functions. Following the convention of the previous analysis [1] and the theoretical paper where the idea first appeared [10], the $S(\alpha,\beta,\gamma,\delta;1)$ notation is used. The parameter $\alpha$ is the Lévy index of stability (or characteristic exponent) that is limited to the domain $0<\alpha\leq 2$. The asymmetry parameter $\beta$ is limited to the domain $-1\leq\beta\leq 1$. The scale parameter $\gamma$ is nonnegative and the location parameter $\delta$ can be any real number with $-\infty<\delta<\infty$. The book by P. J. Nolan [14] details exhaustively the $\alpha=1$ special case and the ubiquitous nature of Lévy-stable source distributions. It provides illustrations of the stable densities in the $S(\alpha,\beta,\gamma,\delta;1)$ parameterization.
@@
The correlation function is thus based on the assumption of the simplest case of univariate and symmetric ($\beta=0$) stable distributions. The scale parameter $\gamma$ is replaced by the physical parameter $R$ [10]. In the standardized notation of Nolan [14], this corresponds to the $(\alpha,\beta{=}0,\gamma{=}R/2^{\frac{1}{\alpha}},\delta;1)$ convention. The Fourier-transformed source-density distribution has a simple form, $f(q)=\exp(iq\delta-\frac{1}{2}|qR|^{\alpha})$ and its modulus square leads to the simple form of the BEC function of Eq. (8). Note that the correlation function does not depend on $\delta$.